If x is equal to 2km what would 4x^2-1 equal

Found 2 solutions by MathLover1, math_tutor2020:
Answer by MathLover1(20849)
Answer by math_tutor2020(3816)


If km refers to kilometers, then,
x = 2
4x^2-1 = 4(2)^2-1
= 4*(4)-1
= 16-1
= 15

Summary:
x = 2 kilometers leads to 4x^2-1 = 15 kilometers

-------------------------------

If km in 2km doesn't refer to kilometers, and instead the k and m are separate variables, then:
x = 2km
4x^2-1 = 4(2km)^2-1
= 4(2^2k^2m^2)-1
= 4(4k^2m^2)-1
= 16k^2m^2-1

Summary:
x = 2km leads to 4x^2-1 = 16k^2m^2-1
if the "km" doesn't refer to kilometers but instead two different variables.